Suppose you just received a shipment of nine televisions. Two of the televisions are defective. If two televisions are randomly​ selected, compute the probability that both televisions work. What is the probability at least one of the two televisions does not​ work?
1
Expert's answer
2020-07-01T18:32:44-0400

we are given that total number of Tvs is 9 and the defective tvs out of 9 is 2 thus we have 7 Tvs working normally. Remember that after first selection there is no replacament.

probability of both TVs working is given by 7/9*6/8= 7/12 = 0.5833

The probability of at least one TV working is given by total probability - probability of both TVs working

probability of at least one not working = 1-7/12 = 5/12 = 0.4167